(1 point) Suppose you want to test the claim the the paired sample data given below come from a population for which the mean difference is
μ
d
=0
μ
d
=
0
.
x
y
60
69
76
64
79
77
57
88
63
67
51
85
73
60


x
60
76
79
57
63
51
73
y
69
64
77
88
67
85
60
Use a
0.01
0.01
significance level to find the following:
(a)
The mean value of the differnces
d
d
for the paired sample data
d
=
d
¯
=
-7.29
(b)
The standard deviation of the differences
d
d
for the paired sample data
s
d
=
s
d
=

(c)
The t test statistic
t=
t
=

(d)
The positive critical value
t=
t
=

(e)
The negative critical value
t=
t
=

(f)
Does the test statistic fall in the critical region?


A. No

B. Yes

(g)
Construct a
99
99
% conficence interval for the population mean of all differences
x−y
x
y
.
<
μ
d
<
<
μ
d
<
Answer #1

(Please round the values to the required number of decimal places.)

(a) The mean value of the differences = -7.29
(b) The standard deviation of the differences = 18.97116
(c) t = -1.0161
(d) The positive critical value =

(e) The negative critical value =

(f) No, it does not.
(g) 99% confidence interval = (-33.86954, 19.29811)